3. (20 points) Streiber Publishing Company, an all-equity firm, generates perpetual earnings before interest and taxes of $2.5 million per year. Its after-tax, all-equity discount rte is 20%. The companys tax rate is 34%. a. What is the value of Streiber Publishing? b. If it adjusts its capital structure to include $600,000 of debt, what is the value of the firm? c. Explain any difference in your answers. d. What assumptions are you making when you are valuing Streiber?
Answer #1

Part a:
We can value the company Streiber Publishing by using the formula below:
EBIT (1 - Tax rate ) / Discount rate
Note: When all the cash flows are perpetual, the cash flow to investors is same as net income after tax which is equal to EBIT (1 - Tax rate ).

Value of the firm= $2500000 (1 – 0.34) / 0.20 = $2500000 (0.66) / 0.20 = $1650000 / 0.20 = $8250000
Thus, $8250000 is the value of the company when there is no debt in its capital structure.

Part b:
When the company includes $600,000 of debt in its capital structure, its value is calculated as:
Value of the company= VL = Vu + Btc
tc is the tax rate
Vu = Present value of an unlevered firm (without debt)
B = kdD/k0=market value of debt (bonds)
Kd is the cost of debt (interest rate).
kdD refers to interest rate multiplied by the debt amount. Thus kdD is the perpetual stream of risk-free payments to debt holders.
k0 = risk-free rate
So, we can say that value of a levered firm=Value of the unlevered firm + Present value of the tax shield provided by debt.
Value of a levered firm= $8250000 + (0.34) ($600000) = $8250000 + $204000 = $8454000


Part c:
The difference in the answers is due to presence of debts. This debt created a tax shield for the company.

Part d:
We have assumed that all the cash flows are perpetual without any growth.